OpenStudy (anonymous):

To buy both green and blue bikes would cost $1500. To buy the green and red bikes would cost $750. To buy al three bikes would cost $2000. How much does each bike cost?

OpenStudy (anonymous):

Let cost of green bike be G, blue bike be B and red bike be R G+B = 1500 ---(1) G+ R= 750 ---(2) G+B+R = 2000. --(3) Take (3) minus (1) G+B+R - (G+B) = 2000-1500 R= 500 since R= 500, from (2) G+R= 750 G+ 500= 750 G= 250 Since G= 250, from (1) G+B = 1500 250 +B = 1500 B= 1250

OpenStudy (hoblos):

red:500 green:250 blue:1250

Directrix (directrix):

Let B = Blue, R, Red, and G, Green B = 1250 R = 500 G = 250 G + B = 1500 G + R = 750 G + R + B = 2000 750 + B = 2000 B = 1250 G + 1250 = 1500 G = 250 250 + R = 750 R = 500